# Anton D. Utyushev,a Vadim I. Zakomirnyi,b Alexey A. Shcherbakov,a Ilia L. Rasskazovc and Alexander Morozd

### v1

@@
[Uncaptioned image] Intuitive understanding of extinction of small particles in absorbing and active host media within the MLWA
Anton D. Utyushev,a Vadim I. Zakomirnyi,b Alexey A. Shcherbakov,a Ilia L. Rasskazovc and Alexander Morozd
[Uncaptioned image] In an absorbing or an active host medium characterized by a complex refractive index $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$, our previously developed modified dipole long-wave approximation (MLWA) is shown to essentially overly with the exact Mie theory results for spherical nanoparticle with radius $a\lesssim 25$ nm ($a\lesssim 20$ nm) in the case of Ag and Au (Al and Mg) nanoparticles. The agreement for Au and Ag (Al and Mg) nanoparticles, slightly better in the case of Au than Ag, continues to be acceptable up to $a\sim 50$ nm ($a\sim 40$ nm), and can be used, at least qualitatively, up to $a\sim 70$ nm ($a\sim 50$ nm) correspondingly. A first order analytic perturbation theory (PT) in a normalized extinction coefficient, $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$, around a nonabsorbing host is developed within the dipole MLWA and its properties are investigated. It is shown that, in a suitable parameter range, the PT can reliably capture the effect of host absorption or gain on the extinction efficiency of various plasmonic nanoparticles.
footnotetext: a School of Physics and Engineering, ITMO University, 197101, Saint-Petersburg, Russiafootnotetext: b Beckman Institute for Advanced Science and Technology, University of Illinois at Urbana-Champaign, Urbana, IL 61801, USAfootnotetext: c SunDensity Inc., Rochester, NY 14604, USA footnotetext: d Wave-scattering.com footnotetext:  Corresponding author: ad.utyushev@gmail.comfootnotetext: † Electronic Supplementary Information (ESI) available: See DOI:
@@ 1 Introduction
Electromagnetic scattering in an absorbing host characterized by a complex refractive index $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$ (where $n_{2}^{\prime}$ and $n_{2}^{\prime\prime}$ are real) has more than fifty years old history. The traditional scattering theory neglects the host dissipation and gain 1, 2, because those cases imply either vanishing or infinite scattering wave at the spatial infinity. Once wave number $k=k^{\prime}+{\rm i}k^{\prime\prime}$ is a complex number, conventional expressions for cross sections cannot be straightforwardly extended for $k^{\prime\prime}\neq 0$, because the expressions yield cross sections as complex quantities. Not surprisingly, the history of scattering in an absorbing host is filled in with a number of controversies 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19. This is probably why in classical textbooks it is only fleetingly mentioned in Section 12.1.3 of ref. 2. Already the definition of an incident intensity is not straightforward, as the field incident on a particle is different at different points on the particle as a result of an absorbing medium 4, 5. In an absorbing or an active unbounded host, the imaginary part $k^{\prime\prime}$ of the complex wave number $k=k^{\prime}+ik^{\prime\prime}=2\pi n_{2}/\lambda_{0}$, where $\lambda_{0}$ is the vacuum wavelength, is nonzero, $k^{\prime\prime}\neq 0$. A great deal of effort was required to arrive at suitable definitions of cross sections for $k^{\prime\prime}\neq 0$.
@@
In contrast to the conventional case of $k^{\prime\prime}=0$, two sets of cross sections are commonly used: inherent and apparent. The former is obtained by performing surface integrals of corresponding Poynting vectors over the particle surface. The approach was developed a long time ago 4, 5, but currently accepted expressions for the inherent cross-sections were not presented until 1999 7, 8, 9, 10. The history of apparent cross sections, which are operationally defined in the far field, began with the optical theorem of Bohren and Gilra 6. Here, too, it took a long time, until 2007, to arrive at definitive expressions for the apparent extinction cross-section cross section, $C_{\rm ext}$, and the so-called apparent scattering cross section, $C_{\rm sca}$ 11, 12, 13, 14. A curious feature of the apparent extinction cross-section in an absorbing host is that it can be negative 20, which is neither an artifact of numerical simulations nor it violates any physical law. The apparent extinction cross-section quantifies the difference in the readings of a forward-scattering detector taken with and without the particle. If the surrounding medium is absorbing, the presence of the particle can in fact make the detector signal stronger, thereby implying a negative extinction cross-section. There is no violation of the energy conservation law, since in this case the extinction cross-section is not used to quantify the energy budget of a finite volume encompassing the particle 20. Another curiosity of apparent cross-sections in an absorbing host is that an intrinsic definition of an apparent absorption cross-sections, $C_{\rm abs}$, is still missing. The difficulty lies in that the very presence of a particle necessarily modifies the near field around the particle. The latter may be the cause of an additional absorption in the host medium outside the particle compared to what is happening in the absence of the particle 8, 9. The critical point is that because the field is disturbed by the particle, additional absorption may be realized in the medium external to the particle 21.
First-principles far-field computations based on the general Lorenz–Mie theory showed that increasing absorption in the host medium broadens and suppresses plasmon resonances in the apparent extinction and effective scattering efficiency factors and suppresses resonance features in the phase function 14. At the same time, using inherent cross-sections it was shown that a surrounding lossy medium can narrow the plasmonic resonances, as well as increase their amplitude dramatically 15, 17. Given this recent revival of interest in the role of the host dissipation or gain on extinction cross sections with conflicting predictions of it either enhancing or reducing the localized surface plasmonic resonance (LSPR) of small particles 14, 15, 17, 18, 19, the focus of present work is on the apparent extinction cross-section $C_{\rm ext}$ in the so-called modified long-wave approximation (MLWA) 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19. The MLWA 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19, which can be viewed as a next-order approximation beyond the Rayleigh limit, is known to overcome a number of severe deficiencies of the Rayleigh limit (see eqn (2) below) and, at least for nonabsorbing hosts, be surprisingly precise 27, 31, 33, 34, 17, 19.
@@
Fig.  1: Schematic representation of the problem under consideration. A spherical particle with radius $a$ and refractive index $n_{1}$ is embedded in an absorbing or gain medium with $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$, where $n_{2}^{\prime}$ and $n_{2}^{\prime\prime}$ are real. Imaginary refractive index of the host $n_{2}^{\prime\prime}>0$ for an absorbing medium and $n_{2}^{\prime\prime}<0$ for a gain medium, and $n_{2}^{\prime\prime}=0$ for a transparent medium. Normalized extinction coefficient $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$ is introduced to quantify the magnitude of the host dissipation or gain.
@@
The outline of our contribution is as follows. Section 2 first recalls in its Subsection 2.1 the expression for distance-independent apparent extinction cross section $C_{\rm ext}$ in the framework of the Lorenz–Mie theory 13. In Subsection 2.2 our earlier developed MLWA 33, 34 is summarized. In Subsection 2.3, still within the framework of MLWA, an analytic first-order perturbation theory in a normalized extinction coefficient, $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$, representing the magnitude of the host dissipation, is developed. In Section 3, performance of the above approximations involving small plasmonic particles in different absorbing and active hosts is investigated. Discussion of some of the observed features is provided in Section 4. We then conclude with Section 5.
@@ 2.1 Apparent extinction cross section in the framework of the Lorenz–Mie theory
The specific expression for distance-independent apparent extinction cross section $C_{\rm ext}$ in the framework of the Lorenz–Mie theory is 13 is given as an infinite sum over different multipole orders $\ell\geq 1$,
@@
In the optical convention, $T_{p\ell}$ correspond to the familiar Mie’s expansion coefficients $a_{\ell}$ and $b_{\ell}$ (eqns (4.53) of ref. 2), i.e. $T_{E\ell}=-a_{\ell}$ for electric, or transverse magnetic (TM), polarization, and $T_{M\ell}=-b_{\ell}$ for magnetic, or transverse electric (TE), polarization. The physical, or operational, meaning of the distance-independent apparent $C_{\rm ext}$ is that it determines the reading of a polarization-sensitive well-collimated radiometer (WCR) at a sufficiently large distance $r$ from the particle 13,
@@
where $\Omega$ is the area of the objective lens of the WCR, and $I_{\rm inc}$ is the intensity of the incident homogeneous (uniform) plane wave at the center of the particle. Importantly, the distance-independent extinction cross-section cannot be introduced in the context of evaluating the energy budget of an arbitrarily shaped volume containing the scattering particle 13.
@@ 2.2 The dipole MLWA
Fig.  2: Demonstration of different physical mechanisms described within MLWA on the extinction spectra of Ag sphere with $a=40$ nm in the host with $n_{2}^{\prime}=1.33$. Here $\bar{\kappa}=0$ for a transparent host and $\bar{\kappa}=0.05$ in case of a dissipative host. The refractive indices of Ag were taken from ref. 35. (a) Quasi-static approximation; (b) radiative correction taking the effect of retardation with respect to the incident field; (c) the MLWA including additionally the dynamic depolarization term. Obviously all the correction terms of the dipole MLWA given by eqn (3) are necessary to achieve a fairly accurate approximation to the exact Mie theory shown in panel (d) for a comparison. All the above approximations, but the exact Mie theory, assume a constant field inside the sphere.
@@
the dynamic depolarization ($D_{p\ell}\sim x^{2}$) and the radiative reaction ($R_{p\ell}\sim x^{2\ell+1}$ for $p=E$ and $R_{p\ell}\sim x^{2\ell+3}$ for $p=M$) in the functional form 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19
@@
Assuming nonmagnetic media, one has $\upsilon=\mu_{1}/\mu_{2}=1$ for magnetic (TE) polarization ($p=M$), and $\upsilon=\varepsilon:=\varepsilon_{1}/\varepsilon_{2}$ for electric (TM) polarization ($p=E$), where the subscript 1 (2) identifies the relevant quantities of a sphere (host). The functional form of eqn (2) makes it transparent that the usual Rayleigh limit, which amounts to setting $D_{p\ell}(x)=R_{p\ell}(x)\equiv 0$ in the denominator for $\ell=1$ and $p=E$, is recovered for $x,x_{s}\ll 1$. Here $x$ is in general complex size parameter, $x=2\pi a/\lambda$, with $\lambda$ being the wavelength in the host medium, whereas $x_{s}=2\pi n_{2}a/n_{1}\lambda=x/\sqrt{\varepsilon}$. The vanishing of the size-independent $F$ in the denominator yields the usual quasi-static Fröhlich LSPR condition, which determines the quasi-static LSPR frequencies $\omega_{0\ell}$. In what follows, we shall focus on the dipole MLWA which yields 33, 34
@@
which corrects formula eqn (6) of ref. 19. The proof of that $\varepsilon_{E1}$ yields complex zero of the denominator $D$ of $a_{1}$ is relegated to Sec. S2 “Complex pole of the dipole contribution within the dipole MLWA” of ESI. In the limit of small $x$, one can expand the denominator of $\varepsilon_{E1}$ as $\sim 1+3x^{2}/5+2ix^{3}/3$, whereby eqn (4) reduces to the familiar classical Bohren and Huffman result (cf. eqn (12.13) of ref. 2):
@@
Fig.  3: Extinction spectra, $Q_{\rm ext}$, for individual spherical Ag nanoparticles with radii $a=10$, $20$, $30$, $40$, $50$ nm in water-like host ($n_{2}^{\prime}=1.33$) with $\bar{\kappa}=0.001,0.005,0.01,0.05$, $0.1$, as shown in the legend for each plot. Spectra were calculated by eqn (1) of the exact Mie theory (solid red line), the dipole MLWA by eqn (3) (blue dot-dashed line), the Mie theory in nonabsorbing host ($n_{2}^{\prime\prime}=0$; dashed green line), and the first order $\bar{\kappa}$-PT of our eqn (9) (dotted orange line). Because the Mie theory in the nonabsorbing host is independent of $\bar{\kappa}$, its plots are identical at each given column. The refractive indices of Ag were taken from ref. 35.
@@ 2.3 Perturbation theory in a normalized extinction coefficient within the MLWA
A useful parametrization of the relative dielectric function $\varepsilon$ between the sphere and the host, suitable for studying the departure from nonabsorbing host, is (cf. eqn (9) of ref. 19):
@@
in the expression eqn (1) of the apparent cross section, $C_{\rm ext}$, it is in principle possible to provide in a systematic way the results for the extinction efficiency, $Q_{\rm ext}=C_{\rm ext}/\pi a^{2}$, in the first order of $\bar{\kappa}$ for both gain and absorbing media in the dipole approximation. The derivative ${\rm d}a_{1}/{\rm d}\bar{\kappa}$ is determined by eqn (S21) of ESI (Section S5: “The derivative ${\rm d}a_{1}/{\rm d}\bar{\kappa}$”),
@@
is the denominator $D$ in eqn (3) in the limit $\bar{\kappa}\to 0$. One finds eventually
@@
which defines the first order expansion of $Q_{\rm ext}$ in the parameter $\bar{\kappa}$ within the perturbation theory (PT). The first term on the rhs is the dipole MLWA in a nonabsorbing host characterized by $\varepsilon_{t}$. The second term on the rhs is the perturbation correction. In what follows, we will refer to eqn (9) as the $\bar{\kappa}$-PT approximation.

### v2

# Intuitive understanding of extinction of small particles in absorbing and active host media within the MLWA

@@
Intuitive understanding of extinction of small particles in absorbing and active host media within the MLWA
@@
[…]
journal: opticajournalarticletype: Research Article
[…]
{abstract*}
In an absorbing or an active host medium characterized by a complex refractive index $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$, our previously developed modified dipole long-wave approximation (MLWA) is shown to essentially overly with the exact Mie theory results for localized surface plasmon resonance of spherical nanoparticles with radius $a\lesssim 25$ nm ($a\lesssim 20$ nm) in the case of Ag and Au (Al and Mg) nanoparticles. The agreement for Au and Ag (Al and Mg) nanoparticles, slightly better in the case of Au than Ag, continues to be acceptable up to $a\sim 50$ nm ($a\sim 40$ nm), and can be used, at least qualitatively, up to $a\sim 70$ nm ($a\sim 50$ nm) correspondingly. A first order analytic perturbation theory (PT) in a normalized extinction coefficient, $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$, around a nonabsorbing host is developed within the dipole MLWA and its properties are investigated. It is shown that, in a suitable parameter range, the PT can reliably isolate and capture the effect of host absorption or host gain on the overall extinction efficiency of various plasmonic nanoparticles.
@@ 1 Introduction
Electromagnetic scattering in an absorbing host characterized by a complex refractive index $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$ (where $n_{2}^{\prime}$ and $n_{2}^{\prime\prime}$ are real) has more than fifty years old history. The traditional scattering theory neglects the host dissipation and gain [1, 2], because those cases imply either vanishing or infinite scattering wave at the spatial infinity. Once wave number $k=k^{\prime}+{\rm i}k^{\prime\prime}=2\pi n_{2}/\lambda_{0}$, where $\lambda_{0}$ is the vacuum wavelength, is a complex number, i.e. $k^{\prime\prime}\neq 0$, conventional expressions for cross sections cannot be straightforwardly extended for $k^{\prime\prime}\neq 0$, because the expressions yield cross sections as complex quantities. Not surprisingly, the history of scattering in an absorbing host is filled in with a number of controversies [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. This is probably why in classical textbooks it is only fleetingly mentioned in Section 12.1.3 of ref. 2. Already the definition of an incident intensity is not straightforward, as the field incident on a particle is different at different points on the particle as a result of an absorbing medium [4, 5].
@@
In contrast to the conventional case of $k^{\prime\prime}=0$, two sets of cross sections are commonly used: inherent and apparent. The former is obtained by performing surface integrals of corresponding Poynting vectors over the particle surface. The approach was developed a long time ago [4, 5], but currently accepted expressions for the inherent cross-sections were not presented until 1999 [7, 8, 9, 10]. The focus of present work will be on the apparent extinction cross-section $C_{\rm ext}$, which is operationally defined in the far field (see Section 2.1 below). The history of apparent cross sections began with the optical theorem of Bohren and Gilra [6]. Here, too, it took a long time, until 2007, to arrive at definitive expressions for the apparent extinction cross-section cross section, $C_{\rm ext}$ [11, 12, 13, 14]. A curious feature of the apparent extinction cross-section in an absorbing host is that it can be negative [20], which is neither an artifact of numerical simulations nor it violates any physical law. The apparent extinction cross-section quantifies the difference in the readings of a forward-scattering detector taken with and without the particle. If the surrounding medium is absorbing, the presence of the particle can in fact make the detector signal stronger, thereby implying a negative extinction cross-section. There is no violation of the energy conservation law, since in this case the extinction cross-section is not used to quantify the energy budget of a finite volume encompassing the particle [20]. Another curiosity of apparent cross-sections in an absorbing host is that an intrinsic definition of an apparent absorption cross-sections, $C_{\rm abs}$, is still missing. The difficulty lies in that the very presence of a particle necessarily modifies the near field around the particle. The latter may be the cause of an additional absorption in the host medium outside the particle compared to what is happening in the absence of the particle [8, 9]. The critical point is that because the field is disturbed by the particle, additional absorption may be realized in the medium external to the particle [21]. Consequently, unlike $C_{\rm abs}$ and $C_{\rm sca}$, only apparent extinction cross-section $C_{\rm ext}$ can be consistently defined (see Section 2.1 below).
The focus of present work is on the apparent extinction cross-section $C_{\rm ext}$ in the so-called modified long-wave approximation (MLWA) [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19]. Firstly, it will be shown that the apparent extinction cross-section conforms better to physical intuition than the intrinsic extinction cross-section. Secondly, the MLWA [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19], which can be viewed as a next-order approximation beyond the Rayleigh limit, is known to overcome a number of severe deficiencies of the Rayleigh limit (see Eq. (2) below) and, at least for nonabsorbing hosts, be surprisingly precise [27, 31, 33, 34, 17, 19].
@@
Figure 1: Schematic representation of the problem under consideration. A spherical particle with radius $a$ and refractive index $n_{1}$ is embedded in an absorbing or gain medium with $n_{2}=n_{2}^{\prime}+{\rm i}n_{2}^{\prime\prime}$, where $n_{2}^{\prime}$ and $n_{2}^{\prime\prime}$ are real. Imaginary refractive index of the host $n_{2}^{\prime\prime}>0$ for an absorbing medium and $n_{2}^{\prime\prime}<0$ for a gain medium, and $n_{2}^{\prime\prime}=0$ for a transparent medium. Normalized extinction coefficient $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$ is introduced to quantify the magnitude of the host dissipation or gain.
@@
The outline of our contribution is as follows. Section 2 first recalls in its Subsection 2.1 the expression for distance-independent apparent extinction cross section $C_{\rm ext}$ in the framework of the Lorenz–Mie theory [13]. In Subsection 2.2 our earlier developed MLWA [33, 34] is summarized. In Subsection 2.3, still within the framework of MLWA, an analytic first-order perturbation theory (PT) in a normalized extinction coefficient, $\bar{\kappa}=n_{2}^{\prime\prime}/n_{2}^{\prime}$, representing the magnitude of the host dissipation, is developed. A motivation for developing the above PT is to isolate and capture the effect of the host absorption or host gain on the overall extinction, which is not possible within the MLWA (the latter captures only overall extinction). In Section 3, performance of the above approximations in describing localized surface plasmon resonances of spherical plasmonic nanoparticles in different absorbing and active hosts is investigated. Discussion of some of the observed features is provided in Section 4. We then conclude with Section 5.
@@ 2.1 Apparent extinction cross section in the framework of the Lorenz–Mie theory
The specific expression for distance-independent apparent extinction cross section $C_{\rm ext}$ in the framework of the Lorenz–Mie theory is [13] is given as an infinite sum over different multipole orders $\ell\geq 1$ (see Fig. 1),
@@
In the optical convention, $T_{p\ell}$ correspond to the familiar Mie’s expansion coefficients $a_{\ell}$ and $b_{\ell}$ (Eqs. (4.53) of ref. 2), i.e. $T_{E\ell}=-a_{\ell}$ for electric, or transverse magnetic (TM), polarization, and $T_{M\ell}=-b_{\ell}$ for magnetic, or transverse electric (TE), polarization. The physical, or operational, meaning of the distance-independent apparent $C_{\rm ext}$ is that it determines the reading of a polarization-sensitive well-collimated radiometer (WCR) at a sufficiently large distance $r$ from the particle [13],
@@
where $\Omega$ is the area of the objective lens of the WCR, and $I_{\rm inc}$ is the intensity of the incident homogeneous (uniform) plane wave at the center of the particle. Importantly, the distance-independent extinction cross-section cannot be introduced in the context of evaluating the energy budget of an arbitrarily shaped volume containing the scattering particle [13].
@@ 2.2 The dipole MLWA
Figure 2: Demonstration of different physical mechanisms described within MLWA on the extinction spectra of Ag sphere with $a=40$ nm in the host with $n_{2}^{\prime}=1.33$. Here $\bar{\kappa}=0$ for a transparent host and $\bar{\kappa}=0.05$ in case of a dissipative host. The refractive indices of Ag were taken from ref. 35. (a) Quasi-static Rayleigh approximation; (b) radiative correction taking the effect of retardation with respect to the incident field; (c) the MLWA including additionally the dynamic depolarization term. Obviously all the correction terms of the dipole MLWA given by Eq. (3) are necessary to achieve a fairly accurate approximation to the exact Mie theory shown in panel (d) for a comparison. All the above approximations, but the exact Mie theory, assume a constant field inside the sphere.
@@
the dynamic depolarization ($D_{p\ell}\sim x^{2}$) and the radiative reaction ($R_{p\ell}\sim x^{2\ell+1}$ for $p=E$ and $R_{p\ell}\sim x^{2\ell+3}$ for $p=M$) in the functional form [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 17, 19]
@@
Assuming nonmagnetic media, one has $\upsilon=\mu_{1}/\mu_{2}=1$ for magnetic (TE) polarization ($p=M$), and $\upsilon=\varepsilon:=\varepsilon_{1}/\varepsilon_{2}$ for electric (TM) polarization ($p=E$), where the subscript 1 (2) identifies the relevant quantities of a sphere (host).
The respective terms $F_{p\ell}$, $D_{p\ell}$, $R_{p\ell}$ in the functional form of Eq. (2) have well-defined physical origin and meaning (see Fig. 2). This allows for an intuitive understanding of scattering from small particles which, as it will be shown, allows for rather reliable substitution of infinite series expansion in terms of Bessel functions of the conventional Mie solution by a single dipole term. The functional form of Eq. (2) makes it transparent that the usual Rayleigh limit, which amounts to setting $D_{p\ell}(x)=R_{p\ell}(x)\equiv 0$ in the denominator for $\ell=1$ and $p=E$, is recovered for $x,x_{s}\ll 1$. Here $x$ is in general complex size parameter, $x=2\pi a/\lambda$, with $\lambda$ being the wavelength in the host medium, whereas $x_{s}=2\pi n_{2}a/n_{1}\lambda=x/\sqrt{\varepsilon}$. The vanishing of the size-independent $F$ in the denominator yields the usual quasi-static Fröhlich condition which determines the quasi-static frequencies $\omega_{0\ell}$ for the occurrence of a localized surface plasmon resonance (LSPR). In what follows, we shall focus on the dipole MLWA which yields [33, 34]
@@
which corrects formula Eq. (6) of ref. 19. The proof of that $\varepsilon_{E1}$ yields complex zero of the denominator $D$ of $a_{1}$ is relegated to Section S3 in Supplement 1. In the limit of small $x$, one can expand the denominator of $\varepsilon_{E1}$ as $\sim 1+3x^{2}/5+2ix^{3}/3$, whereby Eq. (4) reduces to the familiar classical Bohren and Huffman result (cf. Eq. (12.13) of ref. 2):
@@
Figure 3: Extinction spectra, $Q_{\rm ext}$, for individual spherical Ag nanoparticles with radii $a=10$, $20$, $30$, $40$, $50$ nm in water-like host ($n_{2}^{\prime}=1.33$) with $\bar{\kappa}=0.001$, $0.005$, $0.01$, $0.05$, $0.1$, as shown in the legend for each plot. Spectra were calculated by Eq. (1) of the exact Mie theory (solid red line), the dipole MLWA by Eq. (3) (blue dot-dashed line), the Mie theory in nonabsorbing host ($n_{2}^{\prime\prime}=0$; dashed green line), and the first order $\bar{\kappa}$-PT of our Eq. (9) (dotted orange line). Because the Mie theory in the nonabsorbing host is independent of $\bar{\kappa}$, its plots are identical at each given column. The refractive indices of Ag were taken from ref. 35. Note a broadening and suppression of plasmon resonances with increasing host absorption.
@@ 2.3 Perturbation theory in a normalized extinction coefficient within the MLWA
The motivation for developing the perturbation theory is to isolate and capture the effect of the host absorption or host gain on the overall extinction, which is not possible within the MLWA (the latter captures only overall extinction). To this end, a useful parametrization of the relative dielectric function $\varepsilon$ between the sphere and the host, suitable for studying the departure from nonabsorbing host, is (cf. Eq. (9) of ref. 19):
@@
in the expression Eq. (1) of the apparent cross section, $C_{\rm ext}$, it is in principle possible to provide in a systematic way the results for the extinction efficiency, $Q_{\rm ext}=C_{\rm ext}/\pi a^{2}$, in the first order of $\bar{\kappa}$ for both gain and absorbing media in the dipole approximation. The derivative ${\rm d}a_{1}/{\rm d}\bar{\kappa}$ is determined by Eq. S21 from Section S5 in Supplement 1,
@@
is the denominator $D$ in Eq. (3) in the limit $\bar{\kappa}\to 0$. One finds eventually
@@
which defines the first order expansion of $Q_{\rm ext}$ in the parameter $\bar{\kappa}$ within the perturbation theory (PT). The first term on the rhs is the dipole MLWA in a nonabsorbing host characterized by $\varepsilon_{t}$. The second term on the rhs is the perturbation correction. In what follows, we will refer to Eq. (9) as the $\bar{\kappa}$-PT approximation.